# Exploring Disentangled and Controllable Human Image Synthesis: From End-to-End to Stage-by-Stage

### v1

@@ Abstract
Achieving fine-grained controllability in human image synthesis is a long-standing challenge in computer vision. Existing methods primarily focus on either facial synthesis or near-frontal body generation, with limited ability to simultaneously control key factors such as viewpoint, pose, clothing, and identity in a disentangled manner. In this paper, we introduce a new disentangled and controllable human synthesis task, which explicitly separates and manipulates these four factors within a unified framework. We first develop an end-to-end generative model trained on MVHumanNet for factor disentanglement. However, the domain gap between MVHumanNet and in-the-wild data produce unsatisfacotry results, motivating the exploration of virtual try-on (VTON) dataset as a potential solution. Through experiments, we observe that simply incorporating the VTON dataset as additional data to train the end-to-end model degrades performance, primarily due to the inconsistency in data forms between the two datasets, which disrupts the disentanglement process. To better leverage both datasets, we propose a stage-by-stage framework that decomposes human image generation into three sequential steps: clothed A-pose generation, back-view synthesis, and pose and view control. This structured pipeline enables better dataset utilization at different stages, significantly improving controllability and generalization, especially for in-the-wild scenarios. Extensive experiments demonstrate that our stage-by-stage approach outperforms end-to-end models in both visual fidelity and disentanglement quality, offering a scalable solution for real-world tasks. Additional demos are available on the project page: https://taited.github.io/discohuman-project/.
@@
1 Introduction
Synthesizing human-centric images is crucial for various applications like virtual reality, the metaverse, video games, and digital fashion. Advances in GANs [13, 21, 4] and large-scale facial datasets [29, 20] have enabled realistic face generation [19], with subsequent research focusing on controlling facial appearances and expressions [7, 30, 43], though often limited to frontal views. The introduction of 3D-aware GANs, EG3D [5], brought view controllability, paving the way for unified control over viewpoint, expression, and appearance [1, 26, 38, 27, 2]. Compared with human faces, full-body image generation is more complex due to clothing variations and pose deformations.
Early GAN-based full-body image generation methods [11, 12, 35] supporting pose or appearance editing but restricted to near-frontal views. Following the key idea of EG3D [5], EVA3D [16] also tried to further enable view controllability of human image generation. There are also a specific topic in this area which focus on clothing image synthesis, i.e., virtual try-on [6, 32, 25, 8, 45]. To achieve free-view human image synthesis, disentanglement and controllability have been key research focuses. However, a unified framework addressing all requirements remains challenging, primarily due to the lack of datasets with disentangled factors. The recent release of the multi-view human video dataset, MVHumanNet [40], provides the possibility to take a step forward in this research direction.
@@ 1 Introduction
2 Related Work
2.1 Human Synthesis
In the realm of human synthesis, recent works [9, 49, 46, 12, 11, 35, 17, 16, 39] have advanced the field by focusing on generating high-quality, full-body human images with diverse identities, hairstyles, garments, and poses. StyleGan-Human [12] adopts a data-centric perspective to examine the data factors influencing image generation based on the StyleGAN framework. InsetGAN [11] integrates multiple pre-trained GANs, with each focusing on distinct components. To achieve control over pose, local body part appearance and garment style, HumanGAN [35] is capable of performing all tasks related to global appearance sampling, pose transfer, part and garment transfer, as well as part sampling. EVA3D [16] and Get3DHuman [39] expand the synthesis capabilities beyond 2D, pushing toward 3D human generation with realistic texture and pose control. EVA3D [16] incorporates a compositional human NeRF (Neural Radiance Field) representation, which divides the human body into 16 distinct parts and assigns each part an individual network. Each network models the corresponding local volume, allowing EVA3D to capture detailed textures and pose variations with high accuracy. Get3DHuman [39], lifting StyleGan-Human into 3D, advances full-body human synthesis and enables detailed control over body shape, garment, and even facial expression, making it highly applicable in applications like gaming, virtual reality, and digital fashion.
@@ 2.1 Human Synthesis
2.2 Pose Transfer
With the growing prominence of diffusion models, increasing research attention [31, 3, 14, 42, 18, 37] has shifted toward realistic person image generation using the powerful frameworks. PIDM [3] pioneers the utilization of denoising diffusion models for person image synthesis, showcasing their ability to produce high-quality images through iterative noise removal. Expanding on this, PoCoLD [14] incorporates pose constraints into a latent diffusion model, enabling more flexible and controlled image synthesis in the latent space by modulating the subject’s pose. PCDMs [37] introduces a progressive conditional diffusion framework that refines pose-guided image synthesis step-by-step to enhance pose control. CFLD [31] employs a coarse-to-fine latent diffusion framework to enhance control over pose manipulation and improve the realism of synthesized images by progressively refining details from coarse to fine levels in the latent space. MagicAnimate [42] and Animate Anyone [18] further extend diffusion models’ capabilities by focusing on animating person images, enabling smooth transitions between different poses and movements to produce high-quality dynamic effects for virtual characters. Human4DiT [36] proposes a novel framework for generating 360-degree high-quality, spatio-temporally coherent human videos from a single image, combining diffusion transformers for global correlation modeling to achieve fine-grained control over pose and viewpoint.
2.3 Virtual Try-on
Image-based virtual try-on (VTON) synthesizes realistic images by transferring garment appearances onto their bodies. Recent methods use GANs or diffusion models to improve realism and precision. GAN-based approaches like VITON-HD [6] address high-resolution image misalignment with ALIAS normalization, while HR-VITON [25] integrates warping and segmentation into a unified try-on condition generator to reduce artifacts. However, GAN-based models often produce unnatural deformations in wild scenarios, reducing the fidelity and realism of the output result. In contrast, diffusion models have recently demonstrated strong performance in image generation. OOTDiffusion [41] aligns garment features with bodies using Outfitting Fusion, improving realism and controllability without redundant warping. IDM-VTON [8] employs a visual encoder and parallel UNet for better garment encoding, while DressCode [32] introduces a multi-category dataset and a Pixel-level Semantic-Aware Discriminator (PSAD) to enhance image quality and realism. CAT-DM [45] leverages ControlNet to add control conditions and enhance garment feature extraction.
3 Method
3.1 Overview
Fig. 2 provides an overview of the proposed pipelines for human image synthesis: an end-to-end pipeline and a stage-by-stage pipeline. Both pipelines leverage the proposed DiT-based model, DiscoHuman, which enables disentangled and controllable human synthesis.
The end-to-end pipeline directly generates a complete human image from a set of disentangled inputs, including a face image, clothing images (upper/lower clothing or a dress, and shoes), and a pose map, allowing simultaneous control over face identity, clothing, pose, and viewpoint. In contrast, the stage-by-stage pipeline refines the synthesis process in three stages. It first generates an A-pose human image with identity and clothing control, then estimates the back view, and finally synthesizes a free-view image under a target pose and viewpoint. This decomposition enhances controllability and generalizability. DiscoHuman is adapted for the second and third stage by modifying its input configurations. Architectural details of DiscoHuman are further described in Sec. 3.2.
3.2 DiscoHuman
Figure 4: Qualitative comparison of different methods on MVHumanNet [40] datasets. Note that AA stands for AnimateAnyone and MA stands for MagicAnimate.
@@ 3.2 DiscoHuman
A major challenge in implementing VisualDiT is the variation in both the quantity and semantics of input conditions. For instance, a sample may include two clothing images (upper and lower garments) or just one image (a full-body dress). A naive solution would involve designing separate encoders for different semantic categories, but this leads to unacceptable computational overhead.
To address this, the VisualDiT Encoder adopts a unified processing strategy. Given a set of image conditions $I=\{I_{1},I_{2},...,I_{n}\}$, each image is first passed through a VAE encoder $\mathcal{E}_{v}(\cdot)$ to obtain its latent representation. Meanwhile, the corresponding foreground segmentation masks $P=\{P_{1},P_{2},...,P_{n}\}$ are one-hot encoded to indicate their semantic categories. A semantic embedding function $\Phi_{v}(\cdot)$, implemented as a convolutional layer, is applied to these segmentation maps to ensure compatibility with the image latents.
@@
Figure 5: Qualitative comparison of different methods on sampled results from THuman4.0 [50] (bottom) and AvatarRex [51] (top) datasets. Note that AA stands for AnimateAnyone and MA stands for MagicAnimate.
@@ 3.3 End-to-End (E2E) Pipeline
$I_{\text{e2e}}$ =$=\{I_{\text{face}},I_{\text{upper cloth}},I_{\text{lower cloth}},I_{\text{dress}},I_{\text{shoes}}\},$ (6a)
$P_{\text{e2e}}$ =$=\{P_{\text{face}},P_{\text{upper cloth}},P_{\text{lower cloth}},P_{\text{dress}},P_{\text{shoes}}\}.$ (6b)
The input settings for HumanDiT remain consistent with those described in Sec. 3.2. Since a person’s outfit consists of either separate garments (upper and lower) or a whole-body dress, we ensure consistency by inputting a black image with a semantic map labeled as pure ‘background’ whenever a clothing category is missing. To train the e2e model, the forward process is defined as follows:
@@
3.4 Stage-by-Stage (SBS) Pipeline
@@
AnimateAnyone [18] 70.7821 0.4331 13.2582 0.6750 74.1451 0.2642 14.7458 0.7799 69.2732 0.2157 15.5053 0.8264
MagicAnimate [42] 63.4553 0.4798 9.3189 0.7180 96.6649 0.4844 8.3149 0.6974 102.0733 0.5225 7.4551 0.6911
CFLD [31] 56.4594 0.2624 13.4898 0.7878 80.4250 0.2626 13.7125 0.7765 78.8488 0.2281 14.6376 0.8179
[…]
Table 1: Quantitative evaluation on MVHumanNet [40], THuman 4.0 [50], and AvatarReX dataset [51]. The comparison on THuman 4.0 and AvatarReX is conducted in a zero-shot manner. Cells highlighted in denotes the best and second-best performances.
Figure 6: Qualitive zero-shot inference results of different methods on in-the-wild full-body human images generated by Stable Diffusion.
@@
During training, we observe significant lighting shifts between the front and back camera views in the MVHumanNet dataset. Directly using raw MVHumanNet images introduces color saturation shifts, leading to unrealistic textures in the synthesized back views. To mitigate this issue, we employ the FLUX [24] model to enhance the training image quality, particularly in frontal and back A-pose images. As demonstrated in our ablation study in Sec. 4.3, this enhancement significantly improves synthesis quality.
@@
4 Experiments
4.1 Implementation Details
@@
In the implementation of DiscoHuman, both VisualDiT and HumanDiT utilize the pre-trained DiT model from Stable Diffusion 3 Medium [10], which consists of 24 DiT blocks. To conserve GPU memory, VisualDiT employs only the first 12 blocks. The corresponding block features $c^{n}$ from VisualDiT are duplicated across all blocks to ensure alignment with HumanDiT. The synthesis resolution for the image is set to $512\times 768$. For hyperparameter settings, we use an 8-bit Adam optimizer with a learning rate of $2\times 10^{-5}$ and a batch size of 16. All experiments are conducted on two A100 GPUs for 30k training steps. Due to the page limitation, in terms of the dataset preparation, please refer to the supplementary material.
Method FLUX FID ↓ LPIPS ↓ PSNR ↑ SSIM ↑
Back-View Model 38.5276 0.1852 15.3949 0.8077
Back-View Model 46.0028 0.1723 15.6032 0.8179
Free-View Model 20.8115 0.1711 16.9394 0.8275
Free-View Model 27.2246 0.1790 15.9477 0.8090
Table 2: Ablation study on FLUX enhancement for back-view model training of the stage-by-stage pipeline. We mark the best results with .
Method Training Dataset FID ↓ LPIPS ↓ PSNR ↑ SSIM ↑
E2E MVHumanNet + VTON 25.1648 0.4494 10.7573 0.6606
E2E MVHumanNet 20.7037 0.3747 12.5632 0.7294
[…]
SBS MVHumanNet 20.8115 0.1711 16.9394 0.8275
Table 3: Quantitative results of end-to-end (E2E) vs. stage-by-stage (SBS) training under different dataset combinations. We mark the best and second-best results with .
Figure 7: Qualitive comparison for End-to-End (E2E) and Stage-by-Stage (SBS) results on MVHumanNet [40] dataset.
4.2 Comparison
In this section, we primarily present the results of the Stage-by-Stage pipeline, as it achieves the best performance. Since Stage 1 (Front-View Synthesis) relies on pre-trained models, we focus our comparison on Stage 2 and Stage 3 to fairly evaluate our contributions. These two stages together can be interpreted as a pose transfer framework, making them directly comparable to SOTA pose transfer methods.
We therefore compare our approach with AnimateAnyone [18], MagicAnimate [42], and CFLD [31]. Since the official implementation of AnimateAnyone has not been released, we use a third-party reimplementation***Moore-AnimateAnyone: https://github.com/MooreThreads/Moore-AnimateAnyone.
Qualitative Evaluation
Fig. 4 presents the qualitative results on the MVHumanNet dataset. Our method achieves the best performance across various views and poses. For instance, the first example in the third row of Fig. 4 demonstrates a challenging view and pose transfer. Our model accurately generates the corresponding pose for the given view, whereas other methods fail to synthesize a plausible right leg of the input identity.
In addition to the evaluation on MVHumanNet, we also demonstrate our method’s zero-shot capability on the THuman 4.0 [50] and AvatarReX [51] datasets shown in Fig. 5, since none of the comparison methods were trained on these datasets. Nevertheless, our method performs well, while AnimateAnyone and MagicAnimate exhibit lower pose accuracy. As for CFLD, it fails to faithfully preserve the identity of the input reference human. We also apply our Stage-by-Stage framework to in-the-wild human images. From Fig. 6, our method performs well for pose and view control with consistent identity, while other methods fail to faithfully preserve the factors of the input reference human.
Quantitative Evaluation
Tab. 1 presents the quantitative evaluation results of our method. To assess the quality of the synthesized images, we use FID, LPIPS, PSNR, and SSIM as evaluation metrics. Our method achieves the best overall performance across all datasets and metrics.
Notably, the results on THuman 4.0 and AvatarReX are evaluated in a zero-shot setting, further demonstrating the robustness and generalizability of our stage-by-stage (SBS) pipeline. On THuman 4.0 dataset, AnimateAnyone achieves comparable performance in LPIPS, PSNR, and SSIM; however, its FID score is significantly worse than ours, indicating lower perceptual quality. The results show that our method performs consistently well and remains robust across varying numbers of views when generating free-view images. Additionally, CFLD achieves the second-best performance on MVHumanNet, suggesting that it is well-suited for scenarios with fewer viewpoints. However, as the number of views increases, AnimateAnyone outperforms CFLD and becomes the second-best model. Our method, in contrast, remains robust across different view settings, demonstrating its superior adaptability to view variations.
4.3 Ablation Study
During Stage 2 Back-View Synthesis, we find that directly using a front-view A-pose image to predict its back-view counterpart is highly challenging. This difficulty arises not only from the missing information on the back but also from the inconsistent lighting conditions in the MVHumanNet dataset. As a result, if we explicitly use the synthesized back-view results in Stage 3, we observe notable degradation in performance compared to the training setup, which uses a ground-truth A-pose back-view image as input.
To address this, we conduct an ablation study and find that enhancing the quality of front-back A-pose images significantly improves the performance of both the back-view synthesis model and the free-view model. The quantitative results in Tab. 2 support this finding. The first two rows show that applying FLUX [24] to enhance the training A-pose images leads to an improvement in FID. Although LPIPS, PSNR, and SSIM are only marginally affected, this enhancement proves to be valuable, as demonstrated in the last two rows. When generating free-view images using a back-view synthesized by a model trained on an FLUX-enhanced dataset (indicated by a checkmark in the FLUX column), we observe consistent improvements across all evaluation metrics.
5 Discussion: End-to-End vs. Stage-by-Stage
The quantitative results of the End-to-End (E2E) pipeline and the Stage-by-Stage (SBS) pipeline are reported in Tab. 3. Comparing the second and fourth rows, both models are trained solely on MVHumanNet, SBS demonstrates significant improvements in LPIPS, PSNR, and SSIM, though it is slightly less effective in FID. This suggests that SBS better preserves pairwise consistency, especially in image quality after pose transfer.
To further improve synthesis quality, we incorporate VTON data into training for both E2E and SBS pipelines. However, in the E2E pipeline, we observe a decline in performance after introducing VTON data (as shown in the first row). We hypothesize that this degradation is caused by imbalanced conditioning of E2E model with inconsistent data forms between the two datasets. In E2E, MVHumanNet provides controls over face, clothing, shoes, and various poses/views, whereas VTON mainly consists of frontal-view, close-up photographs that often depict only the upper body or lack facial information. When aligning VTON data to be jointly trained with MVHumanNet, the missing disentangled controls are set to $\varnothing$ (e.g., a black image) to match classifier-free guidance training [15]. This misalignment in disentangled factors leads to the observed degradation.
By adopting the stage-by-stage pipeline, we ensure that only Stage 3 (free-view synthesis) is trained on the combined MVHumanNet and VTON datasets. Since VTON lacks front-back A-pose human images, the inputs for front-back view A-pose humans from VTON are set as $\varnothing$. Thus, VTON contributes only to pose-to-image training. Compared to E2E framework, which entangles multiple factors, Stage 3 in SBS focuses solely on pose and texture conditioning. This targeted optimization allows the SBS pipeline trained on MVHumanNet and VTON to achieve significant improvements. As shown in Fig. 7, SBS outperforms E2E, and the strategy trained with VTON dataset produces textures more consistent with the ground truth.
6 Conclusion
We introduce a new and challenging task in human image synthesis that controls viewpoint, pose, clothing, and identity within a unified framework. While an end-to-end model offers a straightforward approach, our analysis reveals its limitations in generalizability, particularly when trained on datasets with inconsistent conditioning factors. To address this, we propose a stage-by-stage framework, which significantly enhances pose and viewpoint control while improving robustness to in-the-wild scenarios. Our experiments demonstrate that structured factor disentanglement leads to improved synthesis quality, particularly in zero-shot settings. By leveraging dataset specialization in different stages, our method enables more effective utilization of available data, outperforming existing approaches in view and pose control. We hope our findings on end-to-end vs. stage-by-stage synthesis will inspire further research in controllable human generation.
Supplementary Material
Figure 8: More results on the Stage-by-Stage (SBS) pipeline.
7 Datasets Preparation

### v4

@@ Abstract
Achieving fine-grained controllability in human image synthesis is a long-standing challenge in computer vision. Existing methods primarily focus on either facial synthesis or near-frontal body generation, with limited ability to simultaneously control key factors such as viewpoint, pose, clothing, and identity in a disentangled manner. In this paper, we introduce a new disentangled and controllable human synthesis task, which explicitly separates and manipulates these four factors within a unified framework. We first develop an end-to-end generative model trained on MVHumanNet for factor disentanglement. However, the domain gap between MVHumanNet and in-the-wild data produces unsatisfactory results, motivating the exploration of virtual try-on (VTON) dataset as a potential solution. Through experiments, we observe that simply incorporating the VTON dataset as additional data to train the end-to-end model degrades performance, primarily due to the inconsistency in data forms between the two datasets, which disrupts the disentanglement process. To better leverage both datasets, we propose a stage-by-stage framework that decomposes human image generation into three sequential steps: clothed A-pose generation, back-view synthesis, and pose and view control. This structured pipeline enables better dataset utilization at different stages, significantly improving controllability and generalization, especially for in-the-wild scenarios. Extensive experiments demonstrate that our stage-by-stage approach outperforms end-to-end models in both visual fidelity and disentanglement quality, offering a scalable solution for real-world tasks.
@@
I Introduction
Synthesizing human-centric images is crucial for various applications like virtual reality, the metaverse, video games, and digital fashion. Advances in GANs [13, 20] and large-scale facial datasets [29, 19] have enabled realistic face generation [18], with subsequent research focusing on controlling facial appearances and expressions [8, 31, 43], though often limited to frontal views. The introduction of 3D-aware GANs, EG3D [5], brought view controllability, paving the way for unified control over viewpoint, expression, and appearance [2, 26, 38, 27, 3]. Compared with human faces, full-body image generation is more complex due to clothing variations and pose deformations.
Early GAN-based full-body image generation methods [11, 12, 35] supporting pose or appearance editing but restricted to near-frontal views. Following the key idea of EG3D [5], EVA3D [15] also tried to further enable view controllability of human image generation. There are also a specific topic in this area which focus on clothing image synthesis, i.e., virtual try-on [6, 32, 24, 7]. To achieve free-view human image synthesis, disentanglement and controllability have been key research focuses. However, a unified framework addressing all requirements remains challenging, primarily due to the lack of datasets with disentangled factors. The recent release of the multi-view human video dataset, MVHumanNet [40, 25], provides the possibility to take a step forward in this research direction.
@@ I Introduction
II Related Work
II-A Human Synthesis
In the realm of human synthesis, recent works [9, 48, 45, 12, 11, 35, 15, 39] have advanced the field by focusing on generating high-quality, full-body human images with diverse identities, hairstyles, garments, and poses. StyleGan-Human [12] adopts a data-centric perspective to examine the data factors influencing image generation based on the StyleGAN framework. InsetGAN [11] integrates multiple pre-trained GANs, with each focusing on distinct components. To achieve control over pose, local body part appearance and garment style, HumanGAN [35] is capable of performing all tasks related to global appearance sampling, pose transfer, part and garment transfer, as well as part sampling. EVA3D [15] and Get3DHuman [39] expand the synthesis capabilities beyond 2D, pushing toward 3D human generation with realistic texture and pose control. EVA3D [15] incorporates a compositional human NeRF (Neural Radiance Field) representation, which divides the human body into 16 distinct parts and assigns each part an individual network. Each network models the corresponding local volume, allowing EVA3D to capture detailed textures and pose variations with high accuracy. Get3DHuman [39], lifting StyleGan-Human into 3D, advances full-body human synthesis and enables detailed control over body shape, garment, and even facial expression, making it highly applicable in applications like gaming and digital fashion.
II-B Pose Transfer
With the rise of diffusion models, increasing research [30, 4, 42, 16, 37] has focused on realistic person image generation. PIDM [4] first explores denoising diffusion for person synthesis, demonstrating high-quality results via iterative refinement. PCDM [37] adopts a step-by-step conditional framework to enhance pose control. MagicAnimate [42] and AnimateAnyone [16] extend diffusion to person animation, achieving smooth pose transitions and dynamic visual effects. Champ [51] incorporates a 3D human model into latent diffusion for enhanced 3D conditioning. Human4DiT [36] leverages diffusion transformers to synthesize 360° coherent human videos from a single image, enabling fine-grained pose and view control.
II-C Virtual Try-on
Image-based virtual try-on (VTON) synthesizes realistic images by transferring garment appearances onto their bodies. Recent methods use GANs or diffusion models to improve realism and precision. GAN-based approaches like VITON-HD [6] address high-resolution image misalignment with ALIAS normalization, while HR-VITON [24] integrates warping and segmentation into a unified try-on condition generator to reduce artifacts. However, GAN-based models often produce unnatural deformations in wild scenarios, reducing the fidelity and realism of the output result. In contrast, diffusion models have recently demonstrated strong performance in image generation. OOTDiffusion [41] aligns garment features with bodies using Outfitting Fusion, improving realism and controllability without redundant warping. IDM-VTON [7] employs a visual encoder and parallel UNet for better garment encoding, while DressCode [32] introduces a multi-category dataset and a Pixel-level Semantic-Aware Discriminator (PSAD) to enhance image quality.
III Method
III-A Overview
Fig. 2 provides an overview of the proposed pipelines for human image synthesis: an end-to-end pipeline and a stage-by-stage pipeline. Both pipelines leverage the proposed DiT-based model, DiscoHuman, which enables disentangled and controllable human image synthesis.
The end-to-end pipeline directly generates a complete human image from a set of disentangled inputs, including a face image, clothing images (upper/lower clothing or a dress, and shoes), and a pose map, allowing simultaneous control over face identity, clothing, pose, and viewpoint. In contrast, the stage-by-stage pipeline refines the synthesis process in three stages. It first generates an A-pose human image with identity and clothing control, then estimates the back view, and finally synthesizes a free-view image under a target pose and viewpoint. This decomposition enhances controllability and generalizability. DiscoHuman is adapted for the second and third stage by modifying its input configurations. Architectural details of DiscoHuman are further described in Sec. III-B.
III-B DiscoHuman
@@
A key challenge in implementing VisualDiT lies in the variability of both the number and semantics of input conditions. For example, a sample may contain two garment images (e.g., top and bottom) or a single image (e.g., a dress). Designing separate encoders for each semantic category is computationally inefficient. To overcome this, the VisualDiT encoder employs a unified strategy. Given a set of input images $I=\{I_{1},I_{2},...,I_{n}\}$, each is encoded via a VAE encoder $\mathcal{E}_{v}(\cdot)$ to produce latent representations. The corresponding segmentation masks $P=\{P_{1},P_{2},...,P_{n}\}$ are one-hot encoded to indicate semantic categories. A convolutional embedding function $\Phi_{v}(\cdot)$ processes these masks to align them with the image latents.
@@
Figure 4: Qualitative comparison of different methods on MVHumanNet [40]. We compare with AnimateAnyone (AA) [16], MagicAnimate (MA) [42], Champ [51], and GPT4o [17]
Figure 5: Qualitative comparison of different methods under the zero-shot setting on sampled results from AvatarReX [50] (top) and THuman4.0 [49] (bottom). We compare with AnimateAnyone (AA) [16], MagicAnimate (MA) [42], Champ [51], and GPT4o [17].
@@ III-C End-to-End (E2E) Pipeline
$I_{\text{e2e}}$ $=\{I_{\text{face}},I_{\text{upper cloth}},I_{\text{lower cloth}},I_{\text{dress}},I_{\text{shoes}}\},$ (6a)
$P_{\text{e2e}}$ $=\{P_{\text{face}},P_{\text{upper cloth}},P_{\text{lower cloth}},P_{\text{dress}},P_{\text{shoes}}\}.$ (6b)
The input settings for HumanDiT remain consistent with those described in Sec. III-B. Since a person’s outfit consists of either separate garments (upper and lower) or a whole-body dress, we ensure consistency by inputting a black image with a semantic map labeled as pure ‘background’ whenever a clothing category is missing. To train the e2e model, the forward process is defined as follows:
@@
III-D Stage-by-Stage (SBS) Pipeline
@@
TABLE I: Quantitative evaluation on MVHumanNet [40], THuman 4.0 [49], and AvatarReX dataset [50]. The comparison on THuman 4.0 and AvatarReX is conducted in a zero-shot manner. Cells highlighted in denotes the best and second-best performances.
@@
AnimateAnyone [16] 70.7821 0.4331 13.2582 0.6750 74.1451 0.3642 14.7458 0.7799 69.2732 0.2157 15.5053 0.8264
MagicAnimate [42] 73.4553 0.4798 9.3189 0.7180 96.6649 0.4844 8.3149 0.6974 102.0733 0.5225 7.4551 0.6911
Champ [51] 62.8560 0.3994 13.1042 0.8635 93.3094 0.4457 9.5117 0.7015 78.9367 0.3717 10.4287 0.7676
Ours (E2E) 25.1648 0.4494 10.7573 0.6606 69.3216 0.3582 12.7761 0.7646 63.0782 0.2091 15.6723 0.8312
Ours (SBS) 14.4210 0.1622 17.0370 0.8311 58.5134 0.2262 14.7664 0.7870 53.0746 0.1744 16.1607 0.8378
Figure 6: Given a reference person image from the People Snapshot dataset [1] (leftmost in each sample) and a target pose (visualized as a colored SMPLX body map), our method (Stage-by-Stage pipeline) synthesizes the same person performing the new action from a specified novel view. Note that our model is evaluated in a zero-shot setting on this dataset, demonstrating strong generalization in pose transfer and view synthesis.
@@
During training, we observe significant lighting shifts between the front and back camera views in the MVHumanNet dataset. Directly using raw MVHumanNet images introduces color saturation shifts, leading to unrealistic textures in the synthesized back views. To mitigate this issue, we employ the FLUX [23] model to enhance the training image quality, particularly in frontal and back A-pose images. As demonstrated in our ablation study in Sec. IV-E, this enhancement significantly improves synthesis quality.
@@
IV Experiments
IV-A Implementation Details
@@
In the implementation of DiscoHuman, both VisualDiT and HumanDiT utilize the pre-trained DiT model from Stable Diffusion 3 Medium [10], which consists of 24 DiT blocks. To conserve GPU memory, VisualDiT employs only the first 12 blocks. The corresponding block features $c^{n}$ from VisualDiT are duplicated across all blocks to ensure alignment with HumanDiT. The synthesis resolution for the image is set to $512\times 768$. For hyperparameter settings, we use an 8-bit Adam optimizer with a learning rate of $2\times 10^{-5}$ and a batch size of 16. All experiments are conducted on two A100 GPUs for 30k training steps.
Figure 7: View interpolation results achieved by the Stage-by-Stage (SBS) pipeline. Each row shows synthesized images of the same identity and clothing under two poses and multiple viewpoints. Despite being an image generation model, DiscoHuman produces smooth and consistent transitions across views, demonstrating strong continuity and factor disentanglement.
Figure 8: Shape control under varying shape parameters. The leftmost image in each example is the input, and the corresponding pose is shown in the top-left corner of each result.
IV-B Datasets Preparation